OpenStudy (anonymous):

when x is added to the numerator and subtracted for the denominator of the fraction (A/B), where A cannot = B and B cannot = 0, the value of the fraction is changed to (C/D). find X

OpenStudy (anonymous):

\[\frac{A+x}{B-x} = \frac{C}{D}\]

OpenStudy (anonymous):

therefore: \[D(A+x) = C(B-x)\]

OpenStudy (anonymous):

distributive property: \[DA + Dx = CB - Cx\]

OpenStudy (anonymous):

rearrange: \[Dx+Cx = CB - DA\]

OpenStudy (anonymous):

combine like terms: \[(D+C)x = CB - DA\]

OpenStudy (anonymous):

and then divide: \[x = \frac{CB - DA}{D+C}\]

OpenStudy (anonymous):

@MrMoose thanks so much

OpenStudy (anonymous):

You are welcome